OpenStudy (anonymous):

@triciaal Variable y varies directly with x^2, and y = 196 when x = 7. Enter the constant of variation in the quadratic variation equation. y = x^2

OpenStudy (anonymous):

Direct variation for quadratics is: y = kx^2 Plug in what we know: 196 = k(7)^2 196 = 49k Divide 49 to both sides, what's 196 / 49?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, so that is the constant, so we can plug it into our equation: y = 4x^2

OpenStudy (anonymous):

Thank u!

OpenStudy (anonymous):

Oh! I am learning Quadratic Variation, too!